# Effective programming of a photonic processor with complex interferometric structure

### v1

@@ Acknowledgements.
The authors acknowledge support from Russian Science Foundation grant 22-12-00353P (https://rscf.ru/en/project/22-12-00353/); V.K., A.D. and G.G. acknowledge support of the Ministry of Science and Higher Education of the Russian Federation FSME-2025-0004 (PICs fabrication).
@@ Appendix A Fabrication process
The waveguide circuit was written using 515-nanometer laser pulses, which were the second harmonic of the Avesta Antaus ytterbium fiber femtosecond laser system. These pulses had a duration of 280 femtoseconds and were delivered at a repetition rate of 1 megahertz. Each pulse had an energy of 120 nanojoules and a linear polarisation parallel to the writing direction inside the fused silica glass sample. The laser beam was focused using an aspheric lens with a numerical aperture (NA) of approximately 0.55, 15 microns below the surface of the sample. A 150-micron-thick cover glass was placed between the lens and the sample to partially correct for spherical aberrations. A reconfigurable beam expander was used to enlarge the input aperture of the focusing lens.A high-precision AeroTech Fiber Glide 3D air-bearing system was used to move the sample at a speed of 0.2 mm/s during the waveguide fabrication process. The average propagation loss is approximately 0.65 dB/cm at 925 nm, and the coupling loss is approximately 1.5 dB per end face. Additionally, the bending loss was less than 0.1 dB/cm for a bending radius of 60 mm used in the experiment.
@@ Appendix B Optical chip design
Our photonic processor contained an auxiliary part of three cascaded Mach-Zhenders’ interferometers (MZIs) and a target $4\times 4$ interferometer consisting of two $4\times 4$ directional couplers and a phase shift layer in between. The design of the entire optical chip, including the waveguide structure and the electrode pattern, is shown in Fig.  5. The radius of curvature for almost all bends is 60 mm, which is a compromise between the bending loss effect and the overall size of the waveguide structure that can be achieved using our manufacturing process. The inner pair of waveguides in the $4\times 4$ directional couplers have approximately three times larger radius of curvature. The input and output ports are spaced 250 microns apart and are connected to single-mode fiber optic arrays.
@@ Appendix C Preparation part calibration
The preparation part of our optical chip comprising of three cascaded MZIs can be calibrated with the help of the four parallel auxiliary waveguides, which couples the output waveguides of the preparation part of the chip as shown in Fig. 6. Therefore, to calibrate the three MZIs of the preparation part, we injected coherent laser radiation in the second input port of the chips’ first (and main) floor of the waveguides and registered the output radiation from the output ports of the auxiliary waveguides - i.e. a second floor of the waveguide structure of our optical chip.
@@
The optical power coupling between the preparation part of the interferometer and the auxiliary waveguides was deliberately made weak ($\leq 10\%$) so as not to loosen much of the signal injected into the target interferometer. However, the amount of coupled radiation was sufficient to calibrate the MZIs.
@@ Appendix D Simultaneous approximation details
As stated in the main text, the necessary data for calibration of the phase shift layer is obtained by consequently injecting coherent radiation into each input port of the target interferometer and applying current in some range at each heater (at once) while measuring the optical power from all the output ports.
@@
    Current was being applied to the heater $j\in\overline{1,3}$ in range from 0 to $15.6$ mA with the step of $0.12$ mA (resulting in $131$ different current values), while the optical power was being measured from each $m$-th output port of the chip (see Fig. 1b). The measured optical power was then normalized.
@@
    The data then is organized in a way that allows each data point to be accessed using four indices: input port number $k\in\overline{1,4}$, heater number $j\in\overline{1,3}$ at which current $x$ was applied, index $i\in\overline{1,131}$ that enumerates the applied current $x_{i}$ values ($x_{0}=0$ mA and $x_{131}=15.6$ mA), and output port number $m\in\overline{1,3}$ (there is no $m=4$ as the output power was normalized $\sum_{m}P_{m}=1$ and therefore three independent output powers left).
@@
Therefore, the complete experimental data contain $4716$ data points, as $4\text{input ports}\times 3\text{heater}\times 3\text{normed output ports}\times 131\text{current values}=4716$. Each data point $P_{exp}(k,j,i,m)\leq 1$ by construction.
@@
The digital model of chip $U_{sim}$ can predict the normalized output optical power from the target interferometer corresponding to a concrete set of $\{k,j,i,m\}$ indices: $P_{sim}(k,j,i,m)$. The main task of calibrating the phase shifts is to find a digital model of the target interferometer that approximates all experimentally measured data points $P_{sim}(k,j,i,m)\approx P_{exp}(k,j,i,m)$. This task is accomplished by optimization of parameters of the chips’ model $U_{sim}$ to maximize the $R^{2}$ coefficient of determination (loss function) as follows:
@@
where $\overline{P_{exp}}=\dfrac{1}{N}\sum_{k,j,i,m}P_{exp}(k,j,i,m)$ and $N$ is the total number of data points, which equals $N=4716$ in our experiment. Thus, the described optimization procedure is nothing but a curve fitting. Further, we need to determine $30$ real parameters: $9$ in each mode mixing matrix $M_{1,2}$, $3$ bias zero phase shifts $\Phi_{0}$ and $9$ elements of the $3\times 3$ crosstalk matrix $A=\{\alpha_{ij}\}$. As the mode mixing matrices $M_{1,2}$ were parameterized by a universal interferometer consisting of triangular meshes of MZIs, all optimization parameters except for nine $\{\alpha_{ij}\}$ are phase shifts that can be varied in the $0,2\pi]$ range. We have not supposed any apriori information about $M_{1,2}$ form and treat them as two independent generally different unitary matrices parametrized by nine real parameters each ($t_{1-9}$ and $t_{10-18}$), whose initial values for optimization were chosen at random from $0,2\pi]$. The initial values for the bias phase shifts, $Phi_{0}$, are always taken as zero. However, the initial guess for the nine crosstalk matrix elements $\{\alpha_{ij}\}$ should have been chosen more accurately, both because it did not have finite bounds of its’ values as phase shifts parameters $t_{j}$ and $\Phi_{0}$ had a $0,2\pi]$, and because it had an explicit physical meaning of thermal distribution in the sample material that must be governed by the geometry of the heaters and the heat equation. Therefore, a preliminary estimate of (at least) the relative values of the crosstalk matrix elements $\{\alpha_{ij}\}$ is necessary.
@@
To estimate the relations between crosstalk matrix elements $\alpha_{ij}$ we considered a simple system of four equivalent (endless) straight and parallel heaters of width $r_{0}$ that are placed above four equivalent (endless) straight and parallel waveguides spaced at a distance $d$ apart from each other (see Fig. 7). Such a model system, perhaps, roughly describes the heaters in our actual target interferometer which are notably not straight, as shown in Fig. 5b. Nevertheless, this simplified model can still be helpful for estimation.
The physical meaning of the crosstalk matrix $A=\{a_{ij}\}$ is how "strong" the $j$-th heater affects the $i$-th heater (or, more precisely, the $i$-th interferometric arm). For example, if current $x_{1}$ is applied to the first heater (see Fig. 7a) it will cause the phase shift change not only in the first waveguide $\varphi_{1}(x_{1})$, on which it is supposed to affect, but in the second $\varphi_{2}(x_{1})$ and third $\varphi_{3}(x_{1})$ waveguides as well. However, the impact of the first heater on the second waveguide is lower than that on the first one, and the impact of the first heater on the third waveguide is even lower than that on the second one. This is because the phase shift induced by the heater is proportional to the temperature of the material in which the waveguide is embedded, and the temperature decreases with the distance from the thermal energy source.
@@
where $T_{0}$ is the maximum temperature of the heater, and $\beta$ is a constant with the dimension of temperature. For the four waveguides in phase shift layer, three relative phase shifts are sufficient. We assume that all three phase shifts are relative to the fourth waveguide: $\varphi_{i}\equiv\varphi_{i}-\varphi_{4}$ for $i\in[1,3]$, which means $\varphi_{i}$ is the phase shift between the $i$-th waveguide and the fourth waveguide (or $\varphi_{4}\equiv 0$). The phase shift $\varphi_{i}$ arises because of the temperature difference between the $i$-th and fourth waveguides: $\varphi_{i}-\varphi_{4}\sim T_{i}-T_{4}$. Therefore, when $j$-th heater receives some current $x$ it leads to a phase shift $\varphi_{ij}$ at $i$-th waveguide because of the temperature distribution in the material, as follows:
@@
, as shown schematically in Fig. 7. Without concretizing the explicit $T(r)$ dependence, it immediately follows from (14) that in the system with the simple geometry of parallel heaters, $\varphi_{23}\approx 0$, as the temperatures at the fourth and second waveguides are equal when current is applied to the third heater (which is due to the symmetry of the simplified system). Moreover, when the third heater was turned on, the temperature at the fourth waveguide was greater than that at the first, which led to the negativity of $\varphi_{13}<0$. Other qualitative relations between induced phase shift $\varphi_{ij}$ can be excluded from (14) such as: $\varphi_{11}>\varphi_{22}>\varphi_{33}>\varphi_{21}>\varphi_{12}\approx\varphi_{32}>\varphi_{31}>\varphi_{23}\approx 0>\varphi_{13}$, and $|\varphi_{13}|\approx\varphi_{12}$, which all follows from the fact that the temperature decreases with distance from the energy source.
The relations between the phase shifts $\varphi_{ij}$ becomes quantitative after substituting the logarithmic law for the temperature in (14):
@@
where $1.71=\log_{1.5}2$ and $2.71=\log_{1.5}3$. As can be seen from (16) relations between non diagonal elements of $A$ are constant for any system of straight parallel heaters (and waveguides), whereas the diagonal elements of $A$ are functions of a dimensionless parameter $\xi=\dfrac{d}{r_{0}}$.
@@
With (17) as an initial guess for $A$, three zeros as a starting point for $\Phi_{0}$ and $18$ random numbers from $0,2\pi]$ for $M_{1,2}$, we launched optimization, and it successfully converged to a value of coefficient of determinacy $R^{2}=0.9972$ on $4716$ data points. The predicted by the model output powers $P_{sim}(k,j,i,m)$ and experimentally measured data points $P_{exp}(k,j,i,m)$ are plotted in Fig. 8 as a curves and dots, respectively.
@@
As we can see from the optimization experimental crosstalk matrix $A_{exp}$ did not go too far from the initial estimation $A_{estim}$, which demonstrates the goodness of the simplified model used for the crosstalk $A_{estim}$ estimation. Nevertheless, a qualitative difference between $A_{exp}$ and $A_{estim}$ can be observed which is a negative slightly non-zero $\alpha_{23}<0$. However, this value can be legitimately explained by the actual geometry of the heaters and waveguides of our chip (see Fig. 5b). In fact, the distance between the third heater and the fourth waveguide is always slightly smaller (or equal at the centre) than the distance between the third heater and second waveguide, which results in higher temperatures at the fourth waveguide compared to the second, leading to a negative phase shift in the second waveguide relative to the fourth.
No special assumptions on the bias (zero current) phase shift $\Phi_{0}$ were made, and their initial guess values were set to zero $\Phi^{init}_{0}$. The Optimized values for $\Phi^{exp}_{0}$ at a wavelength 925 nm appeared to be $\Phi^{exp}_{0}=\{0.23,0.98,1.02\}$ rad. These values can also be explained by the actual geometry of the waveguides in the phase shift layer, as the first and fourth waveguides have the same length $L_{1,4}$ and the second and third waveguides have the same length, which is clearly smaller than $L_{1,4}$ (see Fig. 5b).
Appendix E Unitaries on chip measurement
Preparation interferometer which was consisted of three cascaded MZIs had the main purpose to serve as a tool on chip for stable and automated unitary transformation $U$ of the target interferometer measurement. Measurement of the unitary matrix of the target interferometers was performed by exploiting coherent laser radiation according to [43].
The task of experimental measuring unitary matrix $U_{ij}=|U_{ij}|e^{i\varphi_{ij}}$ is to determine all its’ modules $|U_{ij}|$ and angles $\varphi_{ij}$. Modules $|U_{ij}|$ of the target unitary can be obtained simply by injecting the coherent laser radiation into each single input port of the target interferometer and then measuring the normalized power distribution from all output ports of the target interferometer:
@@ Appendix E Unitaries on chip measurement
where $P_{ij}$ is the measured optical power from the $i$-th output port of the interferometer, whereas radiation is injected solely into the $j$-th input port. By repeating this experimentally simple procedure $N$ times for each input port of the target interferometer, a matrix of its modules $|U_{ij}|$ can be obtained. However, despite the simplicity of such measurement, there is little subtlety in determining the modules $|U_{ij}|$ of a linear optical interferometer. This arises from the possible presence of inhomogeneous losses at the output ports of the target interferometer, which would lead to the non unity sum of the absolute values squared of the matrix elements $S^{row}_{i}=\sum_{j}|U_{ij}|^{2}\neq 1$ that should sum to $1$ for general unitary matrices. Whereas, all the columns sum to $1$ by construction $S^{col}_{j}=\sum_{i}|U_{ij}|^{2}=1$. This inconsistency can be resolved by adjusting the matrix of squared modules $|U_{ij}|^{2}$ multiplying it from both sides by a two diagonal matrices $D_{1}$ and $D_{2}$ with strictly positive elements, representing losses at input and output of the chip, such that new matrix $|U_{ij}|^{2}_{new}$ satisfies the unitary requirements and all its’ rows and all columns sum to $1$. These matrices $D_{1}$ and $D_{2}$ can be found by a straightforward iterative method called the Sinkhorn–Knopp algorithm [46]. Therefore, this rescaling of the initially measured modules $|U_{ij}|$ should be performed, if needed.
Angles $\varphi_{ij}$ can also be extracted from the optical power measurement by means of registering the interference that requires injecting coherent laser radiation into two input ports.
@@
We assume that coherent radiation is injected into the $l_{1}$ and $l_{2}$ input ports of the target interferometer of the form
@@
where $a$ and $be^{i\theta}$ are the complex amplitudes of the input coherent field such that $a^{2}+b^{2}=1$. With no loss of generality lets assume $a$ and $b$ to be real. Therefore, $\theta$ is their relative phase shift. Let us also introduce a parameter $x$ that can control the value of $\theta$. This parameter plays a crucial role in the angles $\varphi_{ij}$ measuring, as in typical experiment coherent radiation is being splitted in two beams, which are then being injected into the $l_{1}$ and $l_{2}$ input ports of the target interferometer. Then, to one of splitted beams an extra phase shift $\theta(x)$ is being added by varying the experimental parameter $x$ and the output optical power distribution $P(x)$ among output ports of the target interferometer is being measured. The angles of the $l_{2}$ column of the target matrix $\varphi_{il_{2}}$ can be extracted from the measured power from $x$ curves $P_{k}^{l_{1},l_{2}}(x)$. To determine angles in all other columns, usually $l_{2}$ is being varied from $1$ to $N$ except for $l_{2}\neq l_{1}$, while $l_{1}$ is being fixed.
It can be shown that while radiation is being injected into $l_{1}$ and $l_{2}$ input ports, the optical power from the $k$-th output port of the target interferometer as a function of $x$ is governed by:
@@
where $C_{l_{1},l_{2}}$ and $D_{l_{1},l_{2}}$ are some constants (depending on the modulus of $U$). It should be noted that in (21) all angles in $l_{1}$-th column of the $U$ matrix were assumed to be zero ($\varphi_{kl_{1}}\equiv 0$ for each $k\in\overline{1,N}$), as they can be associated with irrelevant $N$ phase shifts $P_{out}(\varphi_{kl_{1}})$ following after the target interferometer $\tilde{U}\equiv P_{out}(\varphi_{kl_{1}})\times U$ [43].
@@
We realized the procedure described above on a chip with the help of a preparation interferometer that can prepare any desired coherent input state to be sent to the target interferometer, including those of the form (20). For angles measurement of the unitary $U$ we used thermo-optical heaters at the output of the preparation part. Therefore, in our case, the current served as the $x$ parameter and $\theta(x)=\theta_{0}+\alpha_{\theta}x^{2}$ leading to the $P_{k}^{l_{1},l_{2}}(x)$ dependence of the form:
@@
Therefore, to determine the angles of the unitary transformation of the target interferometer, we measured all the necessary $P_{k}^{l_{1},l_{2}}(x)$ curves and approximated each using ( 22 ), which returns the values of the corresponding shifts in the sines angles $\gamma_{kl_{2}}$. For a fixed $l_{1}$ and $l_{2}$ each obtained angle $\gamma_{kl_{2}}$ has the extra $\theta^{l_{1},l_{2}}_{0}$ term, which can be eliminated by subtraction the $\varphi_{l_{1}l_{2}}$ angle from each obtained $\gamma_{kl_{2}}$ having $(\tilde{\gamma}_{kl_{2}}=\varphi_{kl_{2}}-\varphi_{l_{1}l_{2}})$ for each $k\in\overline{1,4}$ and each $l_{2}\in\overline{1,4},l_{2}\neq l_{1}$.
Eventually, the procedure of the target unitary $U$ reconstruction results in an experimentally determined matrices of modules $|U_{ij}|$ and angles $\tilde{\gamma}_{ij}$. The unitary matrix $\tilde{U}_{ij}$ of interest is constructed using element-wise matrix multiplication as follows:

### v2

@@ Acknowledgements.
The authors acknowledge support from Russian Science Foundation grant 22-12-00353-P (https://rscf.ru/en/project/22-12-00353/); V.K., A.D. and G.G. acknowledge support of the Ministry of Science and Higher Education of the Russian Federation FSME-2025-0004 (PICs fabrication).
@@ Appendix A Fabrication process
The waveguide circuit was written using 515-nanometer laser pulses, which were the second harmonic of the Avesta Antaus ytterbium fiber femtosecond laser system. These pulses had a duration of 280 femtoseconds and were delivered at a repetition rate of 1 megahertz. Each pulse had an energy of 120 nanojoules and a linear polarization parallel to the writing direction inside the fused silica glass sample. The laser beam was focused using an aspheric lens with a numerical aperture (NA) of approximately 0.55, 15 microns below the surface of the sample. A 150-micrometer thick cover glass was placed between the lens and the sample to partially correct for spherical aberrations. A reconfigurable beam expander was used to enlarge the input aperture of the focus lens. A high-precision AeroTech Fiber Glide 3D air-bearing system was used to move the sample at a speed of 0.2 mm/s during the waveguide fabrication process. The average propagation loss is approximately 0.65 dB/cm at 925 nm, and the coupling loss is approximately 1.5 dB per end face. Additionally, the bending loss was less than 0.1 dB/cm for a bending radius of 60 mm used in the experiment.
@@ Appendix B Optical chip design
Our photonic processor contained an auxiliary part of three cascaded Mach-Zhenders’ interferometers (MZIs) and a target $4\times 4$ interferometer consisting of three thermo optical phase shifter sandwiched between two $4\times 4$ multiport directional couplers. The design of the entire optical chip, including the waveguide structure and the electrode pattern, is shown in Fig.  6. The radius of curvature for almost all bends is 60 mm, which is a compromise between the bending loss effect and the overall size of the waveguide structure that can be achieved using our manufacturing process. The inner pair of waveguides in the $4\times 4$ directional couplers has approximately a three times larger radius of curvature. The input and output ports are spaced 250 micrometers apart and are connected to single-mode fiber optic arrays.
@@ Appendix C Preparation part calibration
The preparation part of our optical chip comprising three cascaded MZIs can be calibrated with the help of the four parallel auxiliary waveguides, which couples the output waveguides of the preparation part of the chip as shown in Fig. 7. Therefore, to calibrate the three MZIs of the preparation part, we injected coherent laser radiation in the second input port of the chips’ upper (and main) plane of the waveguides and registered the output radiation from the output ports of the auxiliary waveguides, i.e. a second (down) plane of the waveguide structure of our optical chip.
@@
The optical power coupling between the preparation part of the interferometer and the auxiliary waveguides was deliberately weakened ($\leq 10\%$) to avoid losing much of the signal injected into the target interferometer. Meanwhile, the amount of coupled radiation was sufficient to calibrate the MZIs.
@@ Appendix D Simultaneous approximation details
As stated in the main text, the necessary data for calibration of the phase shift layer is obtained by consequently injecting coherent radiation into each input port of the target interferometer and sweeping current in some range at each heater while measuring the optical power from all the output ports.
@@
    The current applied to the heater $j\in\overline{1,3}$ was applied in the range of 0 to $15.6$ mA with a step of $0.12$ mA (resulting in $131$ different current values), while the optical power was measured from each $m$-th output port of the chip (see Fig. 1b). The measured optical power was then normalized.
@@
    The data was then organized in a way that allowed each data point to be accessed using four indices: input port number $k\in\overline{1,4}$, heater number $j\in\overline{1,3}$ at which current $x$ was applied, index $i\in\overline{1,131}$ that enumerates the applied current values $x_{i}$ ($x_{0}=0$ mA and $x_{131}=15.6$ mA), and output port number $m\in\overline{1,3}$ (there is no $m=4$ as the output power was normalized $\sum_{m}P_{m}=1$ and therefore there are three independent output powers left).
@@
Therefore, the complete experimental data contain $4716$ data points as $4\text{input ports}\times 3\text{heater}\times 3\text{normed output ports}\times 131\text{current values}=4716$. Each data point $P_{exp}(k,j,i,m)\leq 1$ by construction.
@@
The digital model of the chip $U_{sim}$ can predict the normalized output optical power of the target interferometer corresponding to a concrete set of indices $\{k,j,i,m\}$ indices: $P_{sim}(k,j,i,m)$. The main task of calibrating the phase shifters is to find a digital model of the target interferometer that approximates all experimentally measured data points $P_{sim}(k,j,i,m)\approx P_{exp}(k,j,i,m)$. This task is accomplished by optimizing the parameters of the chips’ model $U_{sim}$ to maximize the $R^{2}$ coefficient of determination (loss function) as follows:
@@
where $\overline{P_{exp}}=\dfrac{1}{N}\sum_{k,j,i,m}P_{exp}(k,j,i,m)$ and $N$ is the total number of data points, which is equal to $N=4716$ in our experiment. Thus, the described optimization procedure is generally a curve fitting. Further, we need to determine $30$ real parameters: $9$ in each mode mixing matrix $M_{1,2}$, $3$ bias zero phase shifts $\Phi_{0}$ and $9$ elements of the $3\times 3$ crosstalk matrix $A=\{\alpha_{ij}\}$. As the mode mixing matrices $M_{1,2}$ were parameterized by a universal interferometer consisting of triangular MZI meshes, all optimization parameters except for nine $\{\alpha_{ij}\}$ are phase shifts that can be varied in the $0,2\pi]$ range. We have not supposed any a priori information about the form of $M_{1,2}$ and treat them as two independent generally different unitary matrices parameterized by nine real parameters each ($t_{1-9}$ and $t_{10-18}$), whose initial values for optimization were chosen at random from $0,2\pi]$. The initial values for the bias phase shifts, $Phi_{0}$, are always taken as zero. However, the initial guess for the nine crosstalk matrix elements $\{\alpha_{ij}\}$ should have been chosen more accurately, both because it did not have finite bounds of its’ values as phase shift parameters $t_{j}$ and $\Phi_{0}$ had a $0,2\pi]$, and because it had an explicit physical meaning of thermal distribution in the sample material that must be governed by the geometry of the heaters and the heat equation. Therefore, a preliminary estimate of (at least) the relative values of the crosstalk matrix elements $\{\alpha_{ij}\}$ is necessary.
@@
To estimate the relations between crosstalk matrix elements $\alpha_{ij}$, we considered a simple system of four equivalent (endless) straight and parallel heaters of width $r_{0}$ that are placed above four equivalent (endless) straight and parallel waveguides spaced at a distance $d$ apart from each other (see Fig. 8). Such a model system, perhaps, roughly describes the heaters in our actual target interferometer which are notably not straight, as shown in Fig. 6b. Nevertheless, this simplified model can still be helpful for estimation.
The physical meaning of the crosstalk matrix $A=\{a_{ij}\}$ is how "strong" the $j$-th heater affects the $i$-th heater (or, more precisely, the $i$-th interferometric arm). For example, if current $x_{1}$ is applied to the first heater (see Fig. 8a) it will cause the phase shift change not only in the first waveguide $\varphi_{1}(x_{1})$, which it is supposed to affect, but in the second $\varphi_{2}(x_{1})$ and third $\varphi_{3}(x_{1})$ waveguides as well. However, the impact of the first heater on the second waveguide is less than that of the first one, and the impact of the first heater on the third waveguide is even lower than that of the second. This is because the phase shift induced by the heater is proportional to the temperature of the material in which the waveguide is embedded, and the temperature decreases with the distance from the thermal energy source.
@@
where $T_{0}$ is the maximum temperature of the heater and $\beta$ is a constant with the dimension of the temperature. For the four waveguides in the phase shift layer, three relative phase shifts are sufficient. We assume that all three phase shifts are relative to the fourth waveguide: $\varphi_{i}\equiv\varphi_{i}-\varphi_{4}$ for $i\in[1,3]$, which means that $\varphi_{i}$ is the phase shift between the $i$-th waveguide and the fourth waveguide (or $\varphi_{4}\equiv 0$). The phase shift $\varphi_{i}$ arises due to the temperature difference between the $i$-th and fourth waveguides: $\varphi_{i}-\varphi_{4}\sim T_{i}-T_{4}$. Therefore, when the $j$-th heater receives some current $x$, it leads to a phase shift $\varphi_{ij}$ in the $i$-th waveguide due to the temperature distribution in the material, as follows:
@@
, as shown schematically in Fig. 8. Without concretizing the explicit dependence of the $T(r)$ it immediately follows from (15) that in the system with the simple geometry of parallel heaters, $\varphi_{23}\approx 0$, as the temperatures at the fourth and second waveguides are equal when current is applied to the third heater (which is due to the symmetry of the simplified system). Moreover, when the third heater was turned on, the temperature in the fourth waveguide was higher than that in the first, which led to the negativeness of $\varphi_{13}<0$. Other qualitative relations between the induced phase shift $\varphi_{ij}$ can be excluded from (15) such as: $\varphi_{11}>\varphi_{22}>\varphi_{33}>\varphi_{21}>\varphi_{12}\approx\varphi_{32}>\varphi_{31}>\varphi_{23}\approx 0>\varphi_{13}$, and $|\varphi_{13}|\approx\varphi_{12}$, which all follow from the fact that the temperature decreases with distance from the energy source.
The relations between the phase shifts $\varphi_{ij}$ become quantitative after substituting the logarithmic law for the temperature in (15):
@@
where $1.71=\log_{1.5}2$ and $2.71=\log_{1.5}3$. As can be seen from (17) the relations between the non diagonal elements of $A$ are constant for any system of straight parallel heaters (and waveguides), whereas the diagonal elements of $A$ are functions of a dimensionless parameter $\xi=\dfrac{d}{r_{0}}$.
@@
With (18) as an initial guess for $A$, three zeros as a starting point for $\Phi_{0}$ and $18$ random numbers from $0,2\pi]$ for $M_{1,2}$, we launched optimization and successfully converged to a value of the coefficient of determinacy $R^{2}=0.9972$ at $4716$ data points. The predicted output powers of the model $P_{sim}(k,j,i,m)$ and the experimentally measured data points $P_{exp}(k,j,i,m)$ are plotted in Fig. 9 as curves and dots, respectively.
@@
As we can see from the optimization crosstalk matrix $A_{exp}$ did not go too far from the initial estimation $A_{estim}$, which demonstrates the goodness of the simplified model used for the crosstalk matrix estimation $A_{estim}$. Nevertheless, a qualitative difference between $A_{exp}$ and $A_{estim}$ can be observed that is slightly negative $\alpha_{23}<0$. However, this value can be legitimately explained by the actual geometry of the heaters and waveguides of our chip (see Fig. 6b). In fact, the distance between the third heater and the fourth waveguide is always slightly smaller (or equal at the center) than the distance between the third heater and the second waveguide, resulting in higher temperatures at the fourth waveguide compared to the second, leading to a negative phase shift in the second waveguide relative to the fourth.
No special assumptions on phase shift bias (zero current) $\Phi_{0}$ were made and their initial guess values were set to zero $\Phi^{init}_{0}$. The Optimized values for $\Phi^{exp}_{0}$ at a wavelength 925 nm appeared to be $\Phi^{exp}_{0}=\{0.23,0.98,1.02\}$ rad. These values can also be explained by the actual geometry of the waveguides in the phase shift layer, as the first and fourth waveguides have the same length $L_{1,4}$ and the second and third waveguides have the same length, which is clearly smaller than $L_{1,4}$ (see Fig. 6b).
Appendix E Unitary transformation of the optical chip measurement
The preparation interferometer, which consisted of three cascaded MZIs, was mainly used as a chip-based instrument for accurate and automated unitary transformation measurement of the target interferometer $U$. Measurement of the unitary matrix of the target interferometers was performed by exploiting coherent laser radiation according to [43].
The task of the experimental measurement of the unitary matrix $U_{ij}=|U_{ij}|e^{i\varphi_{ij}}$ is to determine all its’ modules $|U_{ij}|$ and angles $\varphi_{ij}$. Modules $|U_{ij}|$ of the target unitary can be obtained simply by injecting coherent laser radiation into each single input port of the target interferometer and then measuring the normalized power distribution $P_{ij}$ from all output ports of the target interferometer:
@@ Appendix E Unitary transformation of the optical chip measurement
where $P_{ij}$ is the measured optical power from the $i$-th output port of the interferometer, whereas radiation is injected only into the $j$-th input port. Repetition of this experimentally simple procedure $N$ times for each input port of the target interferometer, a matrix of its modules $|U_{ij}|$ can be obtained. However, despite the simplicity of this measurement, there is little subtlety in determining the modules $|U_{ij}|$ of a linear optical interferometer. This arises from the possible presence of inhomogeneous losses at the output ports of the target interferometer, which would lead to the non uniform sum of the absolute values squared of the matrix elements $S^{row}_{i}=\sum_{j}|U_{ij}|^{2}\neq 1$ that should sum to $1$ for general unitary matrices. Whereas, all columns sum to $1$ by construction $S^{col}_{j}=\sum_{i}|U_{ij}|^{2}=1$. This inconsistency can be resolved by adjusting the matrix of squared modules $|U_{ij}|^{2}$ multiplying it from both sides by two diagonal matrices $D_{1}$ and $D_{2}$ with strictly positive elements, representing losses at input and output of the chip, such that the new matrix $|U_{ij}|^{2}_{new}$ satisfies the unitary requirements and all its’ rows and all columns sum to $1$. These matrices $D_{1}$ and $D_{2}$ can be found by a straightforward iterative method called the Sinkhorn–Knopp algorithm [46]. Therefore, this rescaling of the modules initially measured $|U_{ij}|$ should be performed if necessary.
Angles $\varphi_{ij}$ can also be extracted from optical power measurement by registering the interference that requires injecting coherent laser radiation into two input ports.
@@
We assume that coherent radiation is injected into the $l_{1}$ and $l_{2}$ input ports of the target interferometer in the form:
@@
where $a$ and $be^{i\theta}$ are the complex amplitudes of the input coherent field such that $a^{2}+b^{2}=1$. Without loss of generality, assume that $a$ and $b$ are real. Therefore, $\theta$ is their relative phase shift. Let us also introduce a parameter $x$ that can control the value of $\theta$. This parameter plays a crucial role in the measurement of angles $\varphi_{ij}$, as in a typical experiment, coherent radiation split into two beams, which are then injected into the $l_{1}$ and $l_{2}$ input ports of the target interferometer. Then, to one of the split beams an extra phase shift $\theta(x)$ is added by varying the experimental parameter $x$ and the output optical power distribution $P(x)$ among the output ports of the target interferometer is measured. The angles of the $l_{2}$ column of the target matrix $\varphi_{il_{2}}$ can be extracted from the power measured from the $x$ curves $P_{k}^{l_{1},l_{2}}(x)$. To determine angles in all other columns, usually $l_{2}$ is varied from $1$ to $N$ except for $l_{2}\neq l_{1}$, while $l_{1}$ is fixed.
It can be shown that while radiation is injected into the $l_{1}$ and $l_{2}$ input ports, the optical power from the $k$-th output port of the target interferometer as a function of $x$ is governed by:
@@
where $C_{l_{1},l_{2}}$ and $D_{l_{1},l_{2}}$ are some constants (depending on the modulus of $U$). It should be noted that in (22) all angles in the $l_{1}$-th column of the $U$ matrix were assumed to be zero ($\varphi_{kl_{1}}\equiv 0$ for each $k\in\overline{1,N}$), as they can be associated with $N$ irrelevant phase shifts $P_{out}(\varphi_{kl_{1}})$ after the target interferometer $\tilde{U}\equiv P_{out}(\varphi_{kl_{1}})\times U$ [43].
@@
We realized the procedure described above on a chip with the help of a preparation interferometer that can prepare any desired coherent input state to be sent to the target interferometer, including those of the form (21). For the measurement of angles of the unitary $U$, we used thermo-optical heaters at the output of the preparation part. Therefore, in our case, the current served as the $x$ parameter and $\theta(x)=\theta_{0}+\alpha_{\theta}x^{2}$ leading to the $P_{k}^{l_{1},l_{2}}(x)$ dependence of the form:
@@
Therefore, to determine the angles of the unitary transformation of the target interferometer, we measured all the necessary $P_{k}^{l_{1},l_{2}}(x)$ curves and approximated each using ( 23 ), which returns the values of the corresponding shifts in the sine angles $\gamma_{kl_{2}}$. For fixed $l_{1}$ and $l_{2}$ each angle $\gamma_{kl_{2}}$ has the extra term $\theta^{l_{1},l_{2}}_{0}$, which can be eliminated by subtracting angle $\varphi_{l_{1}l_{2}}$ from each $\gamma_{kl_{2}}$ having $(\tilde{\gamma}_{kl_{2}}=\varphi_{kl_{2}}-\varphi_{l_{1}l_{2}})$ for each $k\in\overline{1,4}$ and each $l_{2}\in\overline{1,4},l_{2}\neq l_{1}$.
Eventually, the procedure for the target unitary reconstruction $U$ results in the experimentally determined matrices of modules $|U_{ij}|$ and angles $\tilde{\gamma}_{ij}$. The unitary matrix $\tilde{U}_{ij}$ of interest is constructed using element-wise matrix multiplication as follows: